# Leveraging Collection-Wide Similarities for Unsupervised Document Structure Extraction

### v1

@@ Abstract
Document collections of various domains, e.g., legal, medical, or financial, often share some underlying collection-wide structure, which captures information that can aid both human users and structure-aware models. We propose to identify the typical structure of document within a collection, which requires to capture recurring topics across the collection, while abstracting over arbitrary header paraphrases, and ground each topic to respective document locations. These requirements pose several challenges: headers that mark recurring topics frequently differ in phrasing, certain section headers are unique to individual documents and do not reflect the typical structure, and the order of topics can vary between documents. Subsequently, we develop an unsupervised graph-based method which leverages both inter- and intra-document similarities, to extract the underlying collection-wide structure. Our evaluations on three diverse domains in both English and Hebrew indicate that our method extracts meaningful collection-wide structure, and we hope that future work will leverage our method for multi-document applications and structure-aware models. [^1] Our code and data available at https://github.com/SLAB-NLP/Doc-Structure-Parser
@@
[…]
@@ 1 Introduction
To support both human users and structure-aware models, we propose to identify the typical document structure within the collection (§2). This requires to capture recurring topics across the collection, while abstracting over arbitrary header paraphrases, and ground each topic to respective document locations. For example, in Figure 1 we want to identify “Case Summary”, “Evidence Presented” and “Verdict” as the recurring and prominent topics of a typical document within the collection of legal verdict decisions, as opposed to “Software Development Standards Implications”, which is document specific. This also requires to identify that “Verdict” and “Judgement Desicion” represent the same topic, i.e., abstract over header paraphrases. The colored bounding boxes in Figure 1 represent the grounding unto each document.
Automatically extracting the typical document structure is challenging. While topic boundaries are often loosely defined by explicit headers, it is hard to use them directly to understand collection-level properties, as headers indicating the same information often vary in phrasing, for example “Verdict”, “Judgement Outcome” and “Desicion”. Moreover, some section headers are local to individual document, and do not participate in the global structure. For example, “Software Development Standards Implication” in Figure 1. Finally, while section order provides some signal, it is often inconsistent, and may vary across different documents. The challenge is then align section headers across the collection, while being flexible enough to discard sections that do not reflect a global shared structure.
@@ 3.1 Identifying Document Headers
Since header formatting may vary across domains, we consider their detection a corpus-specific task. We identify header candidates via rule-based heuristics, by leveraging a collection-wide signal that allows us to differentiate between certain patterns which are prominent throughout the collection (e.g., sentence length and capitalization patterns (Gutehrlé and Atanassova, 2022)), while filtering out other document elements which are sometimes stylized as headers (e.g., page numbers, recurring signatures, copyright licensing, etc.).
@@ 3.2 Building a Collection Representation
Dataset Name # docs Domain Language Strict structure Preserve sections order Headers similarity
Form-10K 500 Financial English V V High
CUAD 389 Legal (contracts) English X X Medium
Heb-Verdicts 277 Legal (verdicts) Hebrew X V Medium
@@ 5.1 Experimental Setup
First, our method embeds the document texts using a pre-trained language model. For the two English datasets we use a version of the MPNet language model Song et al. (2020), as it was reported with the highest average performance on sentence embedding and semantic search tasks. [^4] The MPNet version we use  ${}^{,}$ [^5] Performance of different pre-trained English LMs https://www.sbert.net/docs/pretrained_models.html For our Hebrew Verdict Decisions dataset we use a version of AlephBERT (Seker et al., 2021). [^6] The AlephBERT version we use https://huggingface.co/imvladikon/sentence-transformers-alephbert
@@ 5.2 Graph Representation Evaluation
In this evaluation, participants are shown 10 headers, 9 of which are chosen headers from the same community at random, and the last header is chosen at random from outside the community. Participants are then task with identifying the intruder.
@@
We run this evaluation through Amazon Mechanical Turk, with a pool of $12$ participants, where each unique sample was annotated by exactly one participant, for a total of $900$ collected annotations (around $450$ samples for each English dataset). We pay $0.1-0.2$ USD for a single annotation, aiming for an hourly pay of $12$ USD. To ensure quality annotations, we start with a qualification test where participants need to reach accuracy above $0.4$ with confidence above $0.25$ over a few toy examples. In Figure 5 in the Appendix we provide an example of the annotation interface. As for the Hebrew Verdict Decisions, we collect $132$ annotations via a group of $8$ in-house Hebrew-speaking graduate students, who show similar performance to MTurk qualified workers.
@@ 5.4 Qualitative ToC Analysis
Lastly, we suggest a collection-level qualitative evaluation, to explore the quality of the output ToC. Below we provide several key observations derived from our manual mapping between the predicted and gold Form-10K ToC entries, visualized in Figure 3 in the Appendix.
@@ 6 Related Work
A seminal theory of a pragmatic meaning is rhetorical structure theory (RST Mann and Thompson, 1988), that inspired many follow-up work, mostly approaching document structure extraction as a supervised task Arnold et al. (2019); Kang et al. (2022); Xia et al. (2022); Hua and Wang (2022). The need for supervision narrows the application of such tools only to high-resource languages and small number of domains. Our method is unsupervised, does not require labeling, and is applicable to many languages and domains.

### v2

@@ Abstract
Document collections of various domains, e.g., legal, medical, or financial, often share some underlying collection-wide structure, which captures information that can aid both human users and structure-aware models. We propose to identify the typical structure of document within a collection, which requires to capture recurring topics across the collection, while abstracting over arbitrary header paraphrases, and ground each topic to respective document locations. These requirements pose several challenges: headers that mark recurring topics frequently differ in phrasing, certain section headers are unique to individual documents and do not reflect the typical structure, and the order of topics can vary between documents. Subsequently, we develop an unsupervised graph-based method which leverages both inter- and intra-document similarities, to extract the underlying collection-wide structure. Our evaluations on three diverse domains in both English and Hebrew indicate that our method extracts meaningful collection-wide structure, and we hope that future work will leverage our method for multi-document applications and structure-aware models. [^1] Our code and data are available at https://github.com/SLAB-NLP/Doc-Structure-Parser
@@
Gili Lior1,2    Yoav Goldberg1,3    Gabriel Stanovsky1,2 1Allen Institute for AI    2The Hebrew University of Jerusalem    3Bar-Ilan University gili.lior@mail.huji.ac.il
@@ 1 Introduction
To support both human users and structure-aware models, we propose to identify the typical document structure within the collection (§2). This requires to capture recurring topics across the collection, while abstracting over arbitrary header paraphrases, and ground each topic to respective document locations. For example, in Figure 1 we want to identify “Case Summary”, “Evidence Presented” and “Verdict” as the recurring and prominent topics of a typical document within the collection of legal verdict decisions, as opposed to “Software Development Standards Implications”, which is document specific. This also requires to identify that “Verdict” and “Judgement Decision” represent the same topic, i.e., abstract over header paraphrases. The colored bounding boxes in Figure 1 represent the grounding unto each document.
Automatically extracting the typical document structure is challenging. While topic boundaries are often loosely defined by explicit headers, it is hard to use them directly to understand collection-level properties, as headers indicating the same information often vary in phrasing, for example “Verdict”, “Judgement Outcome” and “Decision”. Moreover, some section headers are local to individual document, and do not participate in the global structure. For example, “Software Development Standards Implication” in Figure 1. Finally, while section order provides some signal, it is often inconsistent, and may vary across different documents. The challenge is then align section headers across the collection, while being flexible enough to discard sections that do not reflect a global shared structure.
@@ 3.1 Identifying Document Headers
Since header formatting may vary across domains, we consider their detection a corpus-specific task. We identify header candidates via rule-based heuristics, by leveraging a collection-wide signal that allows us to differentiate between certain patterns which are prominent throughout the collection (e.g., sentence length and capitalization patterns (Gutehrlé and Atanassova, 2022)), while filtering out other document elements which are sometimes styled as headers (e.g., page numbers, recurring signatures, copyright licensing, etc.).
@@ 3.2 Building a Collection Representation
Dataset Name # docs # words Domain Language Strict structure Preserve sections order Headers similarity
Form-10K 500 18M Financial English High
CUAD 389 3M Legal (contracts) English Medium
Heb-Verdicts 277 3.4M Legal (verdicts) Hebrew Medium
@@ 5.1 Experimental Setup
First, our method embeds the document texts using a pre-trained language model. For the two English datasets we use a version of the MPNet language model Song et al. (2020), as it was reported with the highest average performance on sentence embedding and semantic search tasks. [^4] The MPNet version we use  https://huggingface.co/sentence-transformers/all-mpnet-base-v2, [^5] Performance of different pre-trained English LMs https://www.sbert.net/docs/pretrained_models.html For our Hebrew Verdict Decisions dataset we use a version of AlephBERT (Seker et al., 2021). [^6] The AlephBERT version we use https://huggingface.co/imvladikon/sentence-transformers-alephbert
@@ 5.2 Graph Representation Evaluation
In this evaluation, participants are shown 10 headers, 9 of which are chosen headers from the same community at random, and the last header is chosen at random from outside the community. Participants are then tasked with identifying the intruder.
@@
[…]
@@ 5.4 Qualitative ToC Analysis
Lastly, we suggest a collection-level qualitative evaluation, to explore the quality of the output ToC. Below we provide several key observations derived from our manual mapping between the predicted and gold Form-10K ToC entries, visualized in Figure 3.
@@ 6 Related Work
A seminal theory of a pragmatic meaning is rhetorical structure theory (RST Mann and Thompson, 1988), that inspired much follow-up work, mostly approaching document structure extraction as a supervised task Arnold et al. (2019); Kang et al. (2022); Xia et al. (2022); Hua and Wang (2022). The need for supervision narrows the application of such tools only to high-resource languages and small number of domains. Our method is unsupervised, does not require labeling, and is applicable to many languages and domains.
@@ 7 Conclusion
Acknowledgements
We would like to thank the anonymous reviewers for their helpful comments. Members of the SLAB group at the Hebrew university for their productive discussion. This research was supported by grants from Allen Institute for AI, the Israeli Ministry of Science and Technology (grant no. 2088), and the Council for Higher Education.
@@ Appendix A Parameters For Evaluation
Headers detection.
In our implementation we detect headers for each dataset with a different set of heuristics. For example, for the Form-10k dataset, we search for sentences shorter than 10 tokens, with capitalization over > 50% of the words in the sentence. This produces high recall (most headers are caught through this heuristic), but low precision (many of the caught header candidates are not really headers). An example for a scenario where a sentence is wrongly marked as a header candidate is for missing sections, in which the only section content is ‘None’ or ‘Omitted’. Furthermore, some of the header candidates are real headers, but not informative ones, for example ‘Section A’. In such cases, we also drop these sentences from the set of header candidates using regex.